# Policy Mirror Descent with Lookahead

### v2

@@ 5.2 Analysis of Inexact $h$-PMD
As can be directly seen from this formulation, the $\gamma^{h}$ convergence rate from the exact setting generalizes to the inexact setting. Therefore, higher lookahead depths lead to faster convergence rates. Another improvement with relation to Johnson et al. (2023) is that the additive error term has a milder dependence on the effective horizon $1-\gamma$: the term $\frac{2b}{(1-\gamma)(1-\gamma^{h})}$ is smaller for all values of $h\geq 2$ than $\frac{4b}{(1-\gamma)^{2}}$. Larger lookahead depths yield a smaller asymptotic error in terms of the suboptimality gap.
@@
Compared to Theorem 5.1 in Johnson et al. (2023), the dependence on the effective horizon improves by a factor of the order of $1/(1-\gamma)$ for $h>A^{-1}(1-\gamma)^{-1}$ thanks to our Monte Carlo lookahead estimator. Their sample complexity is of order $\tilde{\mathcal{O}}(\frac{1}{\epsilon^{2}(1-\gamma)^{8}})$, whereas for $h>\frac{1}{A(1-\gamma)}$ ours becomes $\tilde{\mathcal{O}}(\frac{1}{\epsilon^{2}(1-\gamma)^{7}})$.
Our sample complexity does not depend on quantities such as the distribution mismatch coefficient used in prior work (Xiao, 2022; Lan, 2023), which may scale with the state space size.
@@ 6.1 Inexact $h$-PMD with Function Approximation
We note that the policy update above can be implicit to avoid computing a policy for every state. It is possible to implement this algorithm using our Monte Carlo planning method in a similar fashion to Lattimore and Szepesvari (2019): use the procedure in section 5.1 to estimate $Q_{h}^{\pi_{k}}(s,a)$ for all state action pairs $(s,a)$ in some set $\mathcal{C}\subset\mathcal{S}\times\mathcal{A}$, and use these estimates as targets for a least squares regression to approximate $Q_{h}^{\pi_{k}}$ by $\Psi\theta_{k}$ at time step $k$. See Appendix C for further details. We conclude this section with a convergence analysis of $h$-PMD with linear function approximation.
@@ 6.2 Convergence Analysis of $h$-PMD with Linear Function Approximation
Our analysis follows the approach of Lattimore and Szepesvari (2019). We make the following standard assumptions.
@@
The proof of Theorem 6.3 can be found in Appendix C.3. Notice that our performance bound does not depend on the state space size like in sec. 5 and depends on the dimension $d$ of the feature space instead. We should notice though that the choice of the set $\mathcal{C}$ is crucial for this result and given by the Kiefer-Wolfowitz theorem (Kiefer and Wolfowitz, 1960) which guarantees that $|\mathcal{C}|=\mathcal{O}(d^{2})$ as discussed in Lattimore and Szepesvari (2019). Moreover, our method for function approximation yields a sample complexity that is instance independent. Existing results for PMD with function approximation Alfano et al. (2023); Yuan et al. (2023) depend on distribution mismatch coefficients which can scale with the size of the state space.
In terms of computational complexity, at each iteration, $h$-PMD uses $O((MA)^{h})$ elementary operations where $M$ is the size of the minibatch of trajectories used, $A$ is the size of the action space and $h$ is the lookahead depth. This computational complexity which scales exponentially in $h$ is inherent to tree search methods. Despite this seemingly prohibitive computational scaling, tree search has been instrumental in practice for achieving state of the art results in some environments (Schrittwieser et al., 2019; Hessel et al., 2022) and modern implementations (notably using GPU-based search on vectorized environments) make tree search much more efficient in practice.
@@ 7 Simulations
We conduct simulations to investigate the effect of the lookahead depth on the convergence rate and illustrate our theoretical findings. We run the $h$-PMD algorithm for different values of $h$ in both exact and inexact settings on the DeepSea environment from DeepMind’s bsuite (Osband et al., 2020) using a grid size of 64 by 64, and a discount factor $\gamma=0.99$. Additional experiments are provided in Appendix D. Our codebase where all our experiments can be replicated is available here: https://gitlab.com/kimon.protopapa/pmd-lookahead.
@@ 8 Related Work
Policy Mirror Descent and Policy Gradient Methods. Motivated by their empirical success (Schulman et al., 2015, 2017), the analysis of PG methods has recently attracted a lot of attention (Agarwal et al., 2021; Bhandari and Russo, 2021; Khodadadian et al., 2022; Bhandari and Russo, 2024; Xiao, 2022; Lan, 2023; Johnson et al., 2023). Among these works, a line of research focused on the analysis of PMD (Xiao, 2022; Lan, 2023; Li and Lan, 2023; Li et al., 2023; Johnson et al., 2023; Alfano et al., 2023) as a flexible algorithmic framework, and its particular instances such as the celebrated natural policy gradient (Khodadadian et al., 2022; Yuan et al., 2023). In particular, for solving unregularized MDPs (which is our main focus), Xiao Xiao (2022) established linear convergence of PMD in the tabular setting with a rate depending on the instance dependent distribution mismatch coefficients, and extended their results to the inexact setting using a generative model. Similar results were established for the particular case of the natural policy gradient algorithm (Bhandari and Russo, 2021; Khodadadian et al., 2022). More recently, Johnson, Pike-Burke, and Rebeschini Johnson et al. (2023) showed a dimension-free rate for PMD with exact policy evaluation using adaptive stepsizes, closing the gap with PI. Notably, they further show that this rate is optimal and their analysis inspired from PI circumvents the use of the performance difference lemma which was prevalent in prior work. We improve over these results using lookahead and we extend our analysis beyond the tabular setting by using linear function approximation. To the best of our knowledge, the use of multi-step greedy policy improvement in PMD and its analysis are novel in the literature.
@@
Policy Iteration with Multiple-Step Policy Improvement. Multi-step greedy policies have been successfully used in applications such as the game of Go and have been studied in a body of works (Efroni et al., 2018a, b; Winnicki and Srikant, 2023; Efroni et al., 2020; Tomar et al., 2020). In particular, multi-step greedy policy improvement has been studied in conjunction with PI in Efroni et al. (2018a); Winnicki and Srikant (2023). Efroni et al. (Efroni et al., 2018a) introduced $h$-PI which incorporates $h$-lookahead to PI and generalized existing analyses on PI with 1-step greedy policy improvement to $h$-step greedy policies. However, their work requires access to an $h$-greedy policy oracle, and the $h$-PI algorithm is unstable in the stochastic setting even for $h=1$ because of potentially large policy updates. We address all these issues in the present work with our $h$-PMD algorithm. Winnicki and Srikant Winnicki and Srikant (2023) showed that a first-visit version of a PI scheme using a single sample path for policy evaluation converges to the optimal policy provided that the policy improvement step uses lookahead instead of a $1$-step greedy policy. Unlike our present work, the latter work assumes access to a lookahead greedy policy oracle. More recently, Alegre et al. Alegre et al. (2023) proposed to use lookahead for policy improvement in a transfer learning setting. Extending Generalized Policy Improvement (GPI) Barreto et al. (2020) which identifies a new policy that simultaneously improves over all previously learned policies each solving a specific task, they introduce $h$-GPI which is a multi-step extension of GPI. We rather consider a lookahead version of PMD including policy gradient methods instead of PI. We focus on solving a single task and provide policy optimization guarantees in terms of iteration and sample complexities.
@@
Lookahead-Based Policy Improvement with Tree Search. Lookahead policy improvement is usually implemented via tree search methods (Browne et al., 2012; Shah et al., 2020; Schrittwieser et al., 2019; Kocsis and Szepesvári, 2006b; Hallak et al., 2023; Efroni et al., 2019). Recently, Dalal et al. Dalal et al. (2022) proposed a method combining policy gradient and tree search to reduce the variance of stochastic policy gradients. The method relies on a softmax policy parametrization incorporating a tree expansion. Our general $h$-PMD framework which incorporates a different lookahead policy improvement step also brings together PG and tree search methods. Morimura et al. Morimura et al. (2022) designed a method with a mixture policy of PG and MCTS for non-Markov Decision Processes. Grill et al. Grill et al. (2020a) showed that AlphaZero as well as several other MCTS algorithms compute approximate solutions to a family of regularized policy optimization problems which bear similarities with our $h$-PMD algorithm update rule for $h=1$ (see sec. 3.2 therein).
@@ 9 Conclusion
In this work, we introduced a novel class of PMD algorithms with lookahead inspired by the success of lookahead policies in practice. We have shown an improved $\gamma^{h}$-linear convergence rate depending on the lookahead depth in the exact setting. We proposed a stochastic version of our algorithm enjoying an improved sample complexity. We further extended our results to scale to large state spaces via linear function approximation with a performance bound independent of the state space size. Our paper offers several interesting directions for future research. A possible avenue for future work is to investigate the use of more general function approximators such as neural networks to approximate the lookahead value functions and scale to large state-action spaces. Recent work for PMD along these lines (Alfano et al., 2023) might be a good starting point. Enhancing our algorithm with exploration mechanisms (see e.g. the recent work O’Donoghue (2023)) is an important future direction. Our PMD update rule might offer some advantage in this regard as was recently observed in the literature (Li and Lan, 2023). Constructing more efficient fully online estimators for the lookahead action values using MCTS and designing adaptive lookahead strategies to select the tree search horizon (Rosenberg et al., 2023) are promising avenues for future work. Our algorithm brings together two popular and successful families of RL algorithms: PG methods and tree search methods. Looking forward, we hope our work further stimulates research efforts to investigate practical and efficient methods enjoying the benefits of both classes of methods.
@@ Appendix A Additional Details and Proofs for Section 4: Exact $h$-PMD
Applying the three point descent lemma (see Lemma 6 in Xiao [2022] with $\phi(\pi)=\langle Q_{h}^{\pi_{k}}(s,\cdot),\pi\rangle$ for any $\pi\in\Pi$ and $\mathcal{C}=\Delta(\mathcal{A})$ with the notation therein), we obtain for any policy $\pi\in\Pi$,
@@
As a warmup, we start by reporting a convergence result for the standard PMD algorithm, recently established in Johnson et al. [2023]. Using the contraction properties of the Bellman operators, it can be shown that the suboptimality gap $\|V^{\pi_{k}}-V^{*}\|_{\infty}$ for PI iterates converges to zero at a linear rate with a contraction factor $\gamma$, regardless of the underlying MDP. This instance-independent convergence rate was generalized to PMD in Johnson et al. [2023]. Before stating this result, we report a useful three-point descent lemma to quantify the improvement of the PMD policy resulting from the proximal step compared to an arbitrary policy.
@@
This lemma results from an application of Lemma 6 in Xiao [2022] which is a variant of Lemma 3.2 in Chen and Teboulle [1993]. This same result was also used in Johnson et al. [2023].
Theorem A.6 (Exact PMD, Johnson et al. [2023]).
@@
We provide here a complete proof of this result as a warmup for our upcoming main result. Although equivalent, notice that our proof is more compact compared to the one provided in Johnson et al. [2023] (see section 6 and Lemma A.1 to A.3 therein) thanks to the use of our convenient notations, and further highlights the relationship between PMD and PI through the explicit use of Bellman operators.
@@
About the cost of computing adaptive stepsizes. There is typically only one greedy policy at each iteration, and in the case that there is more than one it is possible to just pick any arbitrary greedy policy and the lower bound condition on the step size will still be satisfied. The cost of computing the step size is typically the cost of computing a Bregman divergence between two policies. We elaborate on this in the following using the discussion in Johnson et al. [2023] p. 7 in ‘Computing the step size’ paragraph which also applies to our setting. We use the step size condition $\eta_{k}\geq\frac{1}{c_{k}}\lVert\min_{\pi\in G_{h}(V^{\pi_{k}})}D_{\phi}(\pi,\pi_{k})\rVert_{\infty}$. We have a minimum in the condition which means we can just pick a single greedy policy $\tilde{\pi}\in G_{h}(V^{\pi_{k}})$ (which can easily be computed given a lookahead value function in the exact setting or its estimate in an inexact setting). As for the maximum over the state space, this is because we use the same step size in all the states. This condition can readily be replaced by a state dependent step size $\eta_{k}(s)\geq\frac{1}{c_{k}}D_{\phi}(\tilde{\pi}(\cdot|s),\pi_{k}(\cdot|s))$ which is enough for our results to hold.
About the condition on the stepsizes. First, notice that PMD has been analyzed using an increasing stepsize in the exact setting (see e.g. Lan [2023], Xiao [2022], Li et al. [2023]) and this is natural as this corresponds to emulating Policy Iteration. In the present work, notice that we do not require the stepsize to go to infinity and the stepsize does not typically go to infinity under our condition. Recall that we only require the stepsize to satisfy the condition: $\eta_{k}\geq\frac{1}{c_{k}}\lVert\min_{\pi\in\mathcal{G}_{h}(V^{\pi_{k}})}D_{\phi}(\pi,\pi_{k})\rVert_{\infty}.$ As the policy gets closer to the optimal one it should also get closer to a greedy policy (the optimal policy is greedy with respect to its value function) and the term containing a Bregman divergence will vanish (since the optimal policy is already greedy with respect to its own value function). This effect balances the decreasing $c_{k}$. Notice that we typically choose $c_{k}=\gamma^{2h(k+1)}$.

### v3

@@ 5.2 Analysis of Inexact $h$-PMD
As can be directly seen from this formulation, the $\gamma^{h}$ convergence rate from the exact setting generalizes to the inexact setting. Therefore, higher lookahead depths lead to faster convergence rates. Another improvement with relation to [20] is that the additive error term has a milder dependence on the effective horizon $1-\gamma$: the term $\frac{2b}{(1-\gamma)(1-\gamma^{h})}$ is smaller for all values of $h\geq 2$ than $\frac{4b}{(1-\gamma)^{2}}$. Larger lookahead depths yield a smaller asymptotic error in terms of the suboptimality gap.
@@
Compared to Theorem 5.1 in [20], the dependence on the effective horizon improves by a factor of the order of $1/(1-\gamma)$ for $h>A^{-1}(1-\gamma)^{-1}$ thanks to our Monte Carlo lookahead estimator. Their sample complexity is of order $\tilde{\mathcal{O}}(\frac{1}{\epsilon^{2}(1-\gamma)^{8}})$, whereas for $h>\frac{1}{A(1-\gamma)}$ ours becomes $\tilde{\mathcal{O}}(\frac{1}{\epsilon^{2}(1-\gamma)^{7}})$.
Our sample complexity does not depend on quantities such as the distribution mismatch coefficient used in prior work [53, 25], which may scale with the state space size.
@@ 6.1 Inexact $h$-PMD with Function Approximation
We note that the policy update above can be implicit to avoid computing a policy for every state. It is possible to implement this algorithm using our Monte Carlo planning method in a similar fashion to [27]: use the procedure in section 5.1 to estimate $Q_{h}^{\pi_{k}}(s,a)$ for all state action pairs $(s,a)$ in some set $\mathcal{C}\subset\mathcal{S}\times\mathcal{A}$, and use these estimates as targets for a least squares regression to approximate $Q_{h}^{\pi_{k}}$ by $\Psi\theta_{k}$ at time step $k$. See Appendix C for further details. We conclude this section with a convergence analysis of $h$-PMD with linear function approximation.
@@ 6.2 Convergence Analysis of $h$-PMD with Linear Function Approximation
Our analysis follows the approach of [27]. We make the following standard assumptions.
@@
The proof of Theorem 6.3 can be found in Appendix C.3. Notice that our performance bound does not depend on the state space size like in sec. 5 and depends on the dimension $d$ of the feature space instead. We should notice though that the choice of the set $\mathcal{C}$ is crucial for this result and given by the Kiefer-Wolfowitz theorem [22] which guarantees that $|\mathcal{C}|=\mathcal{O}(d^{2})$ as discussed in [27]. Moreover, our method for function approximation yields a sample complexity that is instance independent. Existing results for PMD with function approximation [3, 54] depend on distribution mismatch coefficients which can scale with the size of the state space.
In terms of computational complexity, at each iteration, $h$-PMD uses $O((MA)^{h})$ elementary operations where $M$ is the size of the minibatch of trajectories used, $A$ is the size of the action space and $h$ is the lookahead depth. This computational complexity which scales exponentially in $h$ is inherent to tree search methods. Despite this seemingly prohibitive computational scaling, tree search has been instrumental in practice for achieving state of the art results in some environments [40, 19] and modern implementations (notably using GPU-based search on vectorized environments) make tree search much more efficient in practice.
@@ 7 Simulations
We conduct simulations to investigate the effect of the lookahead depth on the convergence rate and illustrate our theoretical findings. We run the $h$-PMD algorithm for different values of $h$ in both exact and inexact settings on the DeepSea environment from DeepMind’s bsuite [34] using a grid size of 64 by 64, and a discount factor $\gamma=0.99$. Additional experiments are provided in Appendix D. Our codebase where all our experiments can be replicated is available here: https://gitlab.com/kimon.protopapa/pmd-lookahead.
@@ 8 Related Work
Policy Mirror Descent and Policy Gradient Methods. Motivated by their empirical success [41, 42], the analysis of PG methods has recently attracted a lot of attention [1, 6, 21, 7, 53, 25, 20]. Among these works, a line of research focused on the analysis of PMD [53, 25, 28, 29, 20, 3] as a flexible algorithmic framework, and its particular instances such as the celebrated natural policy gradient [21, 54]. In particular, for solving unregularized MDPs (which is our main focus), Xiao [53] established linear convergence of PMD in the tabular setting with a rate depending on the instance dependent distribution mismatch coefficients, and extended their results to the inexact setting using a generative model. Similar results were established for the particular case of the natural policy gradient algorithm [6, 21]. More recently, Johnson, Pike-Burke, and Rebeschini [20] showed a dimension-free rate for PMD with exact policy evaluation using adaptive stepsizes, closing the gap with PI. Notably, they further show that this rate is optimal and their analysis inspired from PI circumvents the use of the performance difference lemma which was prevalent in prior work. We improve over these results using lookahead and we extend our analysis beyond the tabular setting by using linear function approximation. To the best of our knowledge, the use of multi-step greedy policy improvement in PMD and its analysis are novel in the literature.
@@
Policy Iteration with Multiple-Step Policy Improvement. Multi-step greedy policies have been successfully used in applications such as the game of Go and have been studied in a body of works [12, 13, 51, 15, 49]. In particular, multi-step greedy policy improvement has been studied in conjunction with PI in [12, 51]. Efroni et al. [12] introduced $h$-PI which incorporates $h$-lookahead to PI and generalized existing analyses on PI with 1-step greedy policy improvement to $h$-step greedy policies. However, their work requires access to an $h$-greedy policy oracle, and the $h$-PI algorithm is unstable in the stochastic setting even for $h=1$ because of potentially large policy updates. We address all these issues in the present work with our $h$-PMD algorithm. Winnicki and Srikant [51] showed that a first-visit version of a PI scheme using a single sample path for policy evaluation converges to the optimal policy provided that the policy improvement step uses lookahead instead of a $1$-step greedy policy. Unlike our present work, the latter work assumes access to a lookahead greedy policy oracle. More recently, Alegre et al. [2] proposed to use lookahead for policy improvement in a transfer learning setting. Extending Generalized Policy Improvement (GPI) [4] which identifies a new policy that simultaneously improves over all previously learned policies each solving a specific task, they introduce $h$-GPI which is a multi-step extension of GPI. We rather consider a lookahead version of PMD including policy gradient methods instead of PI. We focus on solving a single task and provide policy optimization guarantees in terms of iteration and sample complexities.
@@
Lookahead-Based Policy Improvement with Tree Search. Lookahead policy improvement is usually implemented via tree search methods [9, 43, 40, 24, 18, 14]. Recently, Dalal et al. [11] proposed a method combining policy gradient and tree search to reduce the variance of stochastic policy gradients. The method relies on a softmax policy parametrization incorporating a tree expansion. Our general $h$-PMD framework which incorporates a different lookahead policy improvement step also brings together PG and tree search methods. Morimura et al. [31] designed a method with a mixture policy of PG and MCTS for non-Markov Decision Processes. Grill et al. [16] showed that AlphaZero as well as several other MCTS algorithms compute approximate solutions to a family of regularized policy optimization problems which bear similarities with our $h$-PMD algorithm update rule for $h=1$ (see sec. 3.2 therein).
@@ 9 Conclusion
In this work, we introduced a novel class of PMD algorithms with lookahead inspired by the success of lookahead policies in practice. We have shown an improved $\gamma^{h}$-linear convergence rate depending on the lookahead depth in the exact setting. We proposed a stochastic version of our algorithm enjoying an improved sample complexity. We further extended our results to scale to large state spaces via linear function approximation with a performance bound independent of the state space size. Our paper offers several interesting directions for future research. A possible avenue for future work is to investigate the use of more general function approximators such as neural networks to approximate the lookahead value functions and scale to large state-action spaces. Recent work for PMD along these lines [3] might be a good starting point. Enhancing our algorithm with exploration mechanisms (see e.g. the recent work [33]) is an important future direction. Our PMD update rule might offer some advantage in this regard as was recently observed in the literature [28]. Constructing more efficient fully online estimators for the lookahead action values using MCTS and designing adaptive lookahead strategies to select the tree search horizon [38] are promising avenues for future work. Our algorithm brings together two popular and successful families of RL algorithms: PG methods and tree search methods. Looking forward, we hope our work further stimulates research efforts to investigate practical and efficient methods enjoying the benefits of both classes of methods.
@@ Appendix A Additional Details and Proofs for Section 4: Exact $h$-PMD
Applying the three point descent lemma (see Lemma 6 in [Xiao, 2022] with $\phi(\pi)=\langle Q_{h}^{\pi_{k}}(s,\cdot),\pi\rangle$ for any $\pi\in\Pi$ and $\mathcal{C}=\Delta(\mathcal{A})$ with the notation therein), we obtain for any policy $\pi\in\Pi$,
@@
As a warmup, we start by reporting a convergence result for the standard PMD algorithm, recently established in [Johnson et al., 2023]. Using the contraction properties of the Bellman operators, it can be shown that the suboptimality gap $\|V^{\pi_{k}}-V^{*}\|_{\infty}$ for PI iterates converges to zero at a linear rate with a contraction factor $\gamma$, regardless of the underlying MDP. This instance-independent convergence rate was generalized to PMD in [Johnson et al., 2023]. Before stating this result, we report a useful three-point descent lemma to quantify the improvement of the PMD policy resulting from the proximal step compared to an arbitrary policy.
@@
This lemma results from an application of Lemma 6 in [Xiao, 2022] which is a variant of Lemma 3.2 in [Chen and Teboulle, 1993]. This same result was also used in [Johnson et al., 2023].
Theorem A.6 (Exact PMD, [Johnson et al., 2023]).
@@
We provide here a complete proof of this result as a warmup for our upcoming main result. Although equivalent, notice that our proof is more compact compared to the one provided in [Johnson et al., 2023] (see section 6 and Lemma A.1 to A.3 therein) thanks to the use of our convenient notations, and further highlights the relationship between PMD and PI through the explicit use of Bellman operators.
@@
About the cost of computing adaptive stepsizes. There is typically only one greedy policy at each iteration, and in the case that there is more than one it is possible to just pick any arbitrary greedy policy and the lower bound condition on the step size will still be satisfied. The cost of computing the step size is typically the cost of computing a Bregman divergence between two policies. We elaborate on this in the following using the discussion in [Johnson et al., 2023] p. 7 in ‘Computing the step size’ paragraph which also applies to our setting. We use the step size condition $\eta_{k}\geq\frac{1}{c_{k}}\lVert\min_{\pi\in G_{h}(V^{\pi_{k}})}D_{\phi}(\pi,\pi_{k})\rVert_{\infty}$. We have a minimum in the condition which means we can just pick a single greedy policy $\tilde{\pi}\in G_{h}(V^{\pi_{k}})$ (which can easily be computed given a lookahead value function in the exact setting or its estimate in an inexact setting). As for the maximum over the state space, this is because we use the same step size in all the states. This condition can readily be replaced by a state dependent step size $\eta_{k}(s)\geq\frac{1}{c_{k}}D_{\phi}(\tilde{\pi}(\cdot|s),\pi_{k}(\cdot|s))$ which is enough for our results to hold.
About the condition on the stepsizes. First, notice that PMD has been analyzed using an increasing stepsize in the exact setting (see e.g. [Lan, 2023, Xiao, 2022, Li et al., 2023]) and this is natural as this corresponds to emulating Policy Iteration. In the present work, notice that we do not require the stepsize to go to infinity and the stepsize does not typically go to infinity under our condition. Recall that we only require the stepsize to satisfy the condition: $\eta_{k}\geq\frac{1}{c_{k}}\lVert\min_{\pi\in\mathcal{G}_{h}(V^{\pi_{k}})}D_{\phi}(\pi,\pi_{k})\rVert_{\infty}.$ As the policy gets closer to the optimal one it should also get closer to a greedy policy (the optimal policy is greedy with respect to its value function) and the term containing a Bregman divergence will vanish (since the optimal policy is already greedy with respect to its own value function). This effect balances the decreasing $c_{k}$. Notice that we typically choose $c_{k}=\gamma^{2h(k+1)}$.